OpenStudy (anonymous):

HELP PLEASE!!! Find all solutions to the equation. cos2x + 2 cos x + 1 = 0

OpenStudy (mertsj):

Factor it.

OpenStudy (mertsj):

Set each factor equal to 0 and solve.

OpenStudy (anonymous):

cos2x + 2cosx + 1 = 0 2(cosx)^2 - 1 + 2cosx + 1 = 0 2(cosx)^2 + 2cosx = 0 2 ( cosx )(cosx + 1 ) = 0 cosx = 0 or cosx = -1 so, x = π / 2 or x = 3π / 2 or x= π hope this helps

OpenStudy (nathan917):

Try this http://answers.yahoo.com/question/index?qid=20080205153101AAowpnU

OpenStudy (mertsj):

Did you mean cos(2x) or cos^2x?

OpenStudy (mertsj):

\[\cos ^2x+2\cos x+1=0\]

OpenStudy (mertsj):

Is that the problem? Or is it this:

OpenStudy (mertsj):

\[\cos( 2x)+2\cos x+1=0\]

OpenStudy (anonymous):

cos^2x+2cos x+1=0

OpenStudy (anonymous):

Thank you all for your help!

OpenStudy (nathan917):

Np

OpenStudy (mertsj):

Well then: \[\cos ^2x+2\cos x+1=0\]

OpenStudy (mertsj):

\[(\cos x+1)(\cos x+1)=0\]

OpenStudy (mertsj):

\[\cos x+1=0\]

OpenStudy (mertsj):

\[\cos x=-1\]

OpenStudy (mertsj):

\[x=\pi +2\pi n\]